# Chika Maduabuchi, Ericmoore Jossou, Matteo Bucci

### v3

@@ 3.1 Model Architecture
The MSEG-VCUQ architecture is a hybrid model that integrates U-Net with VideoSAM to enhance phase detection (PD) segmentation from high-speed video (HSV) data, as shown in Figure 2. This design leverages U-Net’s strengths in generating initial, detailed segmentation masks, which capture primary liquid-vapor boundaries. These initial masks are refined using VideoSAM’s advanced transformer-based attention mechanisms and feature extraction capabilities. This dual-stage process allows MSEG-VCUQ to address the inherent complexities of boiling phenomena, including overlapping bubbles, diverse fluid modalities, and challenging lighting conditions, achieving pixel-level precision across various experimental datasets.
@@
Figure 2: Illustration of the integrated process within the MSEG-VCUQ architecture. Initially, fine-tuned U-Net models produce segmentation masks tailored to each fluid modality, capturing primary liquid-vapor boundaries. These masks are then paired with their corresponding images and processed by the VideoSAM transformer. The model refines segmentation outputs through its image encoder and mask decoder, leveraging SAM’s pre-trained components to achieve consistent and high-accuracy HSV segmentation across various experimental conditions.
$p(x)=\max_{i,j\in N(x)}X_{ij}$ (2)
$g_{ij}^{l}=\sum_{m,n}W_{mn}^{l}\cdot h_{i-m,j-n}^{(l-1)}$ (3)
Given the task-specific nature of U-Net and the distinct characteristics of PD data for different fluids, the model undergoes transfer learning. Pre-trained weights from cellular segmentation tasks are fine-tuned on fluid-specific data modalities (Argon, Nitrogen, FC-72, Water) to adapt to the unique visual patterns of boiling phenomena. This is achieved by freezing the initial layers of the U-Net and fine-tuning deeper layers. The optimization process during fine-tuning is governed by:
$W^{*},b^{*}=\arg\min_{W,b}L(Y,\hat{Y})$ (4)
$MHSA(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (7)
Here, $Q$, $K$, $V$ are the query, key, and value matrices, respectively, and $d$ is the dimensionality of the embeddings. The output from the transformer is fed into the mask decoder to produce the final refined segmentation.
$L_{Dice}=1-\frac{2\sum_{i}y_{i}\hat{y}_{i}}{\sum_{i}y_{i}+\sum_{i}\hat{y}_{i}}$ (8)
The combined loss function is:
$L=L_{BCE}+L_{Dice}$ (9)
3.2 Data Collection and Processing
The foundation for building a generalizable model for HSV PD segmentation is rooted in assembling an extensive, diverse multimodal dataset that captures the variability of boiling conditions. We curated video frame-mask pairs from experiments conducted across multiple data modalities, each offering unique fluid properties, heat flux conditions, and imaging setups. Table 1 summarizes the experimental conditions, including Saturated Pool Boiling (SPB) scenarios for Argon, Nitrogen, and FC-72 at a heat flux of 120-170 kW/m2, as well as Flow Boiling (FB) for water at 10 bar pressure and a mass flux of 500 kg/m2s, yielding 7,500 frames per modality.
Table 1: High-Speed Video Modalities for Experimental Data Collection
Modality Conditions Heat Flux (kW/m2) Frames
Argon SPB 120 6K
Nitrogen SPB 120 6K
FC-72 SPB 170 6K
Water FB (500 kg/m2s) 3000 7.5K
The raw HSV frames were converted to grayscale and normalized to enhance feature visibility by subtracting reference frames and adjusting contrast, minimizing background noise for improved segmentation performance. Ground truth segmentation masks were developed through a semi-automated pipeline that combined U-Net-based initial segmentation with expert manual refinement using ImageJ software [50] and adaptive thresholding algorithms [51]. These preliminary masks were refined to meet accuracy standards while minimizing manual labor.
This meticulous data collection and processing pipeline ensures the robustness and generalizability of the VideoSAM model, facilitating high-accuracy segmentation across different boiling conditions and combining multimodal datasets, efficient annotation strategies, and rigorous preprocessing guarantees that the resulting model can handle simple and complex fluid behaviors. This framework is a scalable solution for future HSV studies, emphasizing adaptability and precision.
3.3 U-Net CNN
3.3.1 Model Architecture
$H_{\text{cat}}^{l}=[H_{\text{enc}}^{l};G_{\text{dec}}^{l}]$ (10)
3.3.2 Transfer Learning
Transfer learning enhanced U-Net’s adaptability, leveraging a pre-trained model from a biological segmentation dataset. This approach involved freezing the encoder layers, which capture general features such as edges, corners, and textures while fine-tuning the deeper decoder layers to specialize them for the specific PD segmentation task.
Custom U-Net models were trained for each modality since the dataset spans multiple modalities—including argon, nitrogen, FC-72, and high-pressure water. The poor generalization of a universal model across all modalities drove this decision. Each model was fine-tuned using transfer learning on a small set of annotated frames, typically three for training and two for validation.
3.3.3 Feature Transfer and Data Annotation Pipeline
Table 2: Feature Transfer and Annotation Algorithm
Step Action Description
6 Preprocess for Feature Highlighting Normalizing images by subtracting reference frames and adjusting contrast to enhance feature visibility.
3.4 VideoSAM Model
3.4.1 Model Architecture
VideoSAM is built on the SAM architecture, extending its design to address the specific challenges of HSV PD segmentation. This hybrid model leverages CNNs and transformer-based components to enhance segmentation accuracy for PD tasks involving complex bubble dynamics across diverse modalities. The core architecture of VideoSAM integrates a U-Net CNN for initial mask generation with the Vision Transformer (ViT-H) backbone from SAM. This configuration combines the detailed feature extraction of CNNs with the transformers’ long-range dependency-capturing ability.
$x_{\text{min}}=\min\{x|\text{mask}(x,y)>0\}$ (12)
$x_{\text{max}}=\max\{x|\text{mask}(x,y)>0\}$ (13)
$y_{\text{min}}=\min\{y|\text{mask}(x,y)>0\}$ (14)
$y_{\text{max}}=\max\{y|\text{mask}(x,y)>0\}$ (15)
$\text{patch}(i,j)=\text{image}[(i:i+P),(j:j+P)]$ (16)
The Mask Decoder in VideoSAM is a lightweight transformer designed to refine segmentation masks by integrating image features from the ViT-H encoder with prompt embeddings. Unlike traditional convolutional decoders, this transformer-based component utilizes self-attention mechanisms, allowing VideoSAM to dynamically weigh different input parts. This adaptability is crucial for capturing the varied bubble sizes, contours, and complex dynamics inherent in HSV data. The self-attention framework enables the model to focus on critical regions precisely. At the same time, a feed-forward neural network layer introduces non-linearity, enhancing the model’s ability to capture intricate boundary details. The mask decoder is fine-tuned specifically for HSV PD segmentation tasks, ensuring it can effectively handle boiling bubbles’ multi-scale and dynamic nature. During this process, the parameters of the vision encoder and prompt encoder remain frozen, leveraging SAM’s pre-trained feature extraction capabilities. This approach balances computational efficiency with accuracy, enabling VideoSAM to deliver pixel-level precision in real-time or near-real-time applications. It is well-suited for high-speed, dynamic experimental environments.
To further address segmentation challenges, VideoSAM employs a combined loss function that integrates Dice Coefficient Loss with Cross-Entropy Loss. This mixed loss approach ensures that VideoSAM accurately segments both foreground (bubble regions) and background areas, balancing precise boundary delineation with robustness against class imbalance. Overall, VideoSAM capitalizes on the detailed feature extraction of CNNs and SAM’s transformer-based attention mechanisms to deliver a robust and versatile solution for HSV segmentation in boiling heat transfer studies.
Figure 5: Interactive Segmentation Process of VideoSAM.
3.4.2 Model Training and Inference
3.4.3 Experimental Setup
This study comprises three primary experiments designed to evaluate the segmentation performance of VideoSAM across multiple data modalities and to benchmark its effectiveness against other models. These experiments explore the model’s generalization capabilities, performance across diverse conditions, and comparison with custom CNN-based segmentation models.
Experiment 1: Zero-Shot Generalization Across Modalities
The first experiment assessed VideoSAM’s ability to generalize to unseen data modalities, effectively testing its zero-shot learning capacity. Here, VideoSAM was trained solely on HSV PD frames of the Argon data modality. After training, the model was tested on other distinct data modalities, including Nitrogen, FC-72, and Water, which were not encountered during training. This experiment evaluated the model’s ability to segment novel data modalities without additional fine-tuning accurately. The model’s performance was evaluated using standard segmentation metrics, specifically Intersection over Union (IoU) and F1 Score. Visual inspections complemented quantitative assessments, enabling a nuanced understanding of segmentation quality. The results highlighted VideoSAM’s capacity to outperform the baseline SAM across various modalities, showing superior generalization without prior exposure to the new datasets.
Experiment 2: Performance Across Multiple Modalities
The second experiment aimed to validate VideoSAM’s robustness in handling multiple data modalities by training the model on a combination of four datasets—Argon, Nitrogen, FC-72, and Water. This comprehensive training approach exposed VideoSAM to diverse experimental conditions, allowing it to better understand modality-specific segmentation requirements. Following training, the model was evaluated on unseen data from each modality to determine its adaptability to new samples within the trained categories. Consistent performance was expected across all datasets, particularly for complex experimental conditions. IoU and F1 Score metrics were used for quantitative assessment, and the findings demonstrated that VideoSAM maintained high segmentation accuracy across all modalities, effectively capturing the intricate dynamics of each experimental setup. VideoSAM outperformed SAM in all cases, underscoring its robustness and adaptability.
Experiment 3: Comparison with Custom U-Net CNN
3.5 Model Evaluation
3.5.1 Boiling Performance Metrics
$\theta_{dry}=1-\frac{\text{Wet Pixels}}{\text{Total Pixels}}$ (17)
$\rho_{cl}=\frac{\text{Contact Line Length}}{\text{Total Pixels}}$ (18)
Table 3: Comprehensive Algorithm for Boiling Metric Computation
Input Binary mask (binaryMask), where 1 represents dry pixels and 0 represents wet pixels.
1 Read binary mask binaryMask.
9 Return $\theta_{dry}$ and $\rho_{cl}$.
3.5.2 Machine Learning Metrics
Metric Formula
Accuracy $A=\frac{TP+TN}{TP+TN+FP+FN}$
Precision $P=\frac{TP}{TP+FP}$
Recall $R=\frac{TP}{TP+FN}$
Specificity $S=\frac{TN}{TN+FP}$
F1 Score $F_{1}=2\times\frac{P\times R}{P+R}=\frac{2\times TP}{2TP+FP+FN}$
IoU (Jaccard Index) $IoU=\frac{TP}{TP+FP+FN}$
3.6 Uncertainty Quantification
Experimental bubble data were extracted to validate these calculations, and their dimensions (radius, area, and perimeter) were measured to calibrate the error estimates. This calibration aligns the simulations with real-world data, enhancing the reliability of uncertainty estimates. A weighted frequency analysis was applied to further refine error quantification, accounting for the varying occurrence frequencies of bubbles of different sizes, thus providing a comprehensive view of the distribution of errors.
Input Domain length $L$, NumSimulations, N_values, R_values
Output Results with PRE and ME for area and perimeter
1 Initialize simulation parameters
2 for each $N$ in N_values do
3    for each $R$ in R_values do
4     Initialize $A_{\text{theoretical}}$, $P_{\text{theoretical}}$ for $R$
5     for $k=1$ to NumSimulations do
6      Randomly position bubble within domain
7      Calculate Dis, define bub (Dis $<$ $R$)
8      Compute $A_{\text{bub}}$, $P_{\text{bub}}$ using pixel count
9     end for
10     Compute PRE and ME for area, perimeter
11     Store errors in Results
12    end for
13 end for
14 Load and process Results for visualization
15 Extract error matrices from Results
16 Visualize errors (Histograms, Surface plots)
17 Estimate errors for experimental data:
18    Match experimental data with simulation parameters
19    Read off errors for specific experimental conditions
20 Perform weighted frequency analysis on errors
21 Output refined error estimations
@@ 4 Results and Discussions
4.1 U-Net CNN Results
4.1.1 Comparative Analysis of Dry Area Fraction and Contact Line Density
Similarly, the bottom left plot in Figure 7 shows contact line density versus mean heat flux. Both methods indicate an increase in contact line density with rising heat flux, highlighting the intensification of the liquid-vapor interface, a hallmark of enhanced boiling activity. However, the U-Net CNN consistently provides higher contact line density values, particularly at higher flux levels, underscoring its superior sensitivity to intricate details in the boiling interface. This divergence becomes prominent around heat flux steps 8-9 (approximately 140-158 kW/m²), where the boiling process introduces more smaller, closely packed bubbles, challenging thresholding to detect accurately.
@@ 4.1.1 Comparative Analysis of Dry Area Fraction and Contact Line Density
4.1.2 3D Histogram Analysis of Heat Flux vs. Bubble Sizes Distribution
The histogram reveals a comprehensive distribution of bubble sizes across a broad range of heat fluxes for the segmentation method using the U-Net CNN. The distribution is skewed towards smaller bubble sizes at lower heat flux values, with a concentration in the lower bubble size region. As the heat flux increases, the histogram shifts toward larger bubble sizes, suggesting a progression toward producing larger vapor bubbles under higher thermal conditions. This trend is typically associated with intensified boiling activity. The color gradient in the histogram emphasizes the prevalence of medium-sized bubbles, which appear most frequently, underscoring the U-Net CNN’s ability to detect a diverse range of bubble sizes accurately.
In contrast, the thresholding histogram shows a similar trend but with some differences. Although larger bubbles are prevalent at higher heat flux levels, the count of smaller bubbles appears reduced compared to the U-Net CNN results, indicating potential limitations in thresholding for detecting finer bubbles, especially at lower heat fluxes. This results in a less nuanced distribution, particularly in the lower bubble-size regions, which may impact bubble characterization accuracy, as thresholding might overlook smaller bubbles effectively captured by the U-Net CNN.
4.1.3 Comparison of Bubble Size Distribution Using Segmentation and Thresholding
The third column illustrates the thresholding results using color coding for size differentiation. However, this technique highlights larger bubbles more prominently while potentially merging or overlooking smaller bubbles, particularly those nearby. This difference suggests that while thresholding effectively identifies bubbles, it may lack the fine-grained sensitivity of the segmentation approach for capturing smaller or closely situated bubbles.
4.1.4 Ground Truth Benchmark
Standardizing ground truth for fine-tuning the U-Net CNN model and validating its segmentation masks is critical, as it is a benchmark for comparison. This study established ground truth by sending the same frames to five independent users with domain expertise in HSV PD analysis. Each user manually annotated the images, delineating bubbles from the background based on their interpretation. These annotations were then compared for dry area fraction and contact line density to understand potential variations.
The top plot in Figure 9 illustrates the variation in dry area fraction and contact line density as measured by different users and image processing techniques for frames selected from videos 8 and 9. This comparison is vital to gauge the effect of subjective human analysis on the selection of ground truth, especially in image processing contexts where objectivity is often challenging. The bar chart segments data for the two selected frames, showing values recorded by each of the five users (User 1 through User 5) alongside the results from two computational methods: U-Net-based segmentation and thresholding (as implemented by Chavagnat et al. [51]). The low variability among the user-derived values highlights minor perceptual inconsistencies or interpretational differences. However, all user annotations cluster around an average of approximately 0.17 for dry area and 0.05 for contact line density. Despite subtle individual variations, this consistency underscores a general alignment in human perception of these metrics.
4.1.5 Perimeter Visualization for Multimodal Segmentation Accuracy
After fine-tuning, the enhanced U-Net model was compared to expert-verified ground truth perimeters, as shown in the combined visualization. The yellow outlines indicate regions of overlap between the model’s predictions and expert annotations, demonstrating that the fine-tuned U-Net model achieved high accuracy, closely matching human-generated segmentation. This iterative process underscores the adaptability of U-Net CNNs in refining segmentation quality for complex fluids like high-pressure water, highlighting their potential for reliable and detailed analysis in industrial applications.
4.1.6 Statistical Comparison of Segmentation Metrics Across Modalities
In the argon and nitrogen modalities, a high consistency is observed between the U-Net and binarization methods regarding dry area fraction and contact line density. Specifically, nitrogen shows minimal deviation, with dry area fractions from segmentation ranging between 0.46 and 0.48 and contact line densities from 0.06 to 0.07. Similarly, the PDFs and CDFs for nitrogen suggest close alignment, with both segmentation techniques yielding values within a narrow band. Argon exhibits slightly more variation, with the segmented dry area fraction spanning 0.48 to 0.53, and binarization results showing a broader range from 0.51 to 0.56. Box plots further highlight this similarity, with mean values for dry area and contact line density metrics from both segmentation methods agreeing within ±5%, indicating consistent performance across these metrics.
4.1.7 Performance of Segmentation Metrics Across Modalities
Figures 12 to 14 illustrate the U-Net CNN model’s segmentation performance and efficacy in capturing boiling dynamics across various experimental conditions and modalities, including argon, nitrogen, FC-72, and high-pressure water (Case 1 and Case 2). The comparative analysis leverages standard segmentation metrics (accuracy, F1 score, IoU, MCC) alongside specific boiling metrics (dry area fraction and contact line density) to understand model performance under diverse conditions.
Segmentation Metrics Across Modalities (Figure 12)
Across argon, nitrogen, and FC-72, the U-Net CNN segmentation model demonstrates high accuracy, F1-score, and MCC, especially with argon and nitrogen, where metric values are notably consistent. This suggests that the model is effective in these simpler boiling cases, where bubbles are larger and fewer small bubbles are present. The FC-72 data, however, exhibits more variability across these metrics, likely due to the greater presence of smaller bubbles and complex boiling patterns that introduce segmentation challenges. The IoU, sensitive to false positives and negatives, shows the greatest variability, particularly in FC-72. This indicates that while the model effectively identifies true positives, the smaller and more complex bubble patterns in FC-72 pose challenges in achieving precise overlap with the ground truth.
Error Analysis and Statistical Insights (Figures 14)
4.2 Uncertainty Quantification
$\mathrm{ME}=\psi_{\mathrm{theo}}-\overline{\psi}_{\mathrm{disc}}$ (20)
4.2.1 Experimental Data Grounding
4.2.2 Convergence Test of Discretization Errors in Bubble Measurements
5K Iterations:
10K Iterations:
15K Iterations:
20K Iterations:
At this final milestone, perimeter, and area PRE show minimal oscillations. The area PRE is particularly stable, oscillating close to zero, indicating strong convergence. The perimeter PRE also shows minimal variability, supporting the conclusion that the simulation has reached a stable numerical solution.
4.2.3 Impact of Bubble Radius and Grid Resolution on Measurement Accuracy
@@
4.2.4 Effects of Erosion and Dilation on Measurement Accuracy
$W=\frac{\sum_{i=1}^{k}\left(v_{i}\cdot w_{i}\right)}{\sum_{i=1}^{k}w_{i}}$ (21)
Table 6: Uncertainty Table
1 184 -0.5 -1.6 -16.3 -1.0
2 110 0.2 5.6 -1.6 -0.3
3 59 0.03 2.4 2.8 0.9
4 31 0.01 1.8 4.8 2.2
5 11 -0.01 -2.2 5.9 3.5
6 7 0.003 1.6 7.1 6.1
7 3 0.003 1.6 8.0 6.1
8 2 0.006 8.1 8.0 10.1
Figure 17: Difference Between Erosion and Dilation
Table 7 summarizes how these boundary modifications influence uncertainties. The analysis revealed that while the area PRE remained constant, perimeter PRE decreased from 13.8% to 6.6% under dilation. This suggests that dilation smooths boundaries, reducing-edge artifacts that cause perimeter overestimation during erosion. Consequently, dilation results in a closer approximation to the actual bubble perimeter, underscoring the importance of selecting appropriate boundary conditions in digital image processing.
Table 7: Effects of Dilation on the Uncertainty Quantification
Erosion -0.4 13.8 1.2e-4
Dilation -0.4 6.6 7.1e-5
4.2.5 Extended Analysis of Uncertainty Across Modalities
@@
Modality Weighted Avg PRE Area (%) Weighted Avg PRE Perimeter (%) Weighted Avg ME Perimeter ($\mu$m)
@@
Implications of Measurement Uncertainty
4.3 VideoSAM Results
4.3.1 Zero-Shot Generalization Across Modalities
This experiment evaluates the zero-shot performance of VideoSAM on unseen data modalities, including the Nitrogen, FC-72, and Water datasets, with both qualitative and quantitative analyses depicted on the left-handed plots of Figure 18. The segmentation results depicted in the top-left plots of Figure 18 highlight VideoSAM’s impressive ability to generalize across different datasets, especially in complex environments like Nitrogen and FC-72. VideoSAM’s binary masks closely align with the ground truth, maintaining precise contours and boundaries of bubbles, which the SAM model struggles with. The SAM model’s segmentation displays inconsistencies, particularly noticeable in high-density regions where VideoSAM excels at capturing intricate bubble structures. This robustness in VideoSAM’s segmentation highlights its suitability for environments with diverse and irregular bubble formations, as seen in the FC-72 dataset.
4.3.2 Multi-Modal Segmentation Performance Analysis

### v4

@@ 3 Methodology
The MSEG-VCUQ architecture is a hybrid model that integrates U-Net with VideoSAM to enhance phase detection (PD) segmentation from high-speed video (HSV) data, as shown in Figure 2. This design leverages U-Net’s strengths in generating initial, detailed segmentation masks, which capture primary liquid-vapor boundaries. To enhance generalization, these initial masks are refined using VideoSAM’s advanced transformer-based attention mechanisms and feature extraction capabilities. This dual-stage process allows MSEG-VCUQ to address the inherent complexities of boiling phenomena, including overlapping bubbles, diverse fluid modalities, and challenging lighting conditions, achieving pixel-level precision across various experimental datasets. Further architectural details, equations, and illustrations of the workflow are provided in Appendix B.1.
@@
Figure 2: Illustration of the integrated process within the VideoSAM architecture. Initially, fine-tuned U-Net models produce segmentation masks tailored to each fluid modality, capturing primary liquid-vapor boundaries. These masks are then paired with their corresponding images and processed by the VideoSAM transformer. The model refines segmentation outputs through its image encoder and mask decoder, leveraging SAM’s pre-trained components to achieve consistent and high-accuracy HSV segmentation across various experimental conditions.
Data Collection and Processing
To establish a robust and generalizable segmentation framework, we curated a multimodal high-speed video (HSV) dataset that captures a wide range of boiling conditions. The dataset includes multiple fluids and experimental setups, ensuring diversity in boiling regimes, fluid properties, and imaging conditions. A systematic preprocessing pipeline was implemented to enhance data quality and optimize model performance. This pipeline includes grayscale conversion, patchification, and semi-automated annotation of boiling video frames. A detailed explanation of the data acquisition, experimental configurations, and annotation strategies is provided in Appendix B.2.
The experimental dataset comprises video frame-mask pairs collected under both Saturated Pool Boiling (SPB) and Flow Boiling (FB) at 500 kg/m2s conditions. These scenarios encompass fluids with distinct thermal and physical properties recorded at high heat fluxes and pressures. Table 1 summarizes the key experimental conditions, including the boiling modality, heat flux (q), and the number of frames available for each fluid.
Table 1: Summary of Experimental Conditions for Data Collection
Modality Condition q (kW/m2) Frames
Argon SPB 120 6,000
Nitrogen SPB 120 6,000
FC-72 SPB 170 6,000
Water FB 3,000 7,500
This multimodal dataset is the foundation for training and validating segmentation models, ensuring their applicability across different fluids and boiling scenarios.
U-Net CNN
The U-Net model is the backbone for preliminary segmentation, leveraging transfer learning and a feature-rich encoder-decoder architecture. We fine-tune custom U-Net for each fluid modality, with key components such as skip connections and upsampling enhancing its ability to capture bubble structures. Technical details on the U-Net architecture, transfer learning strategy, and the feature transfer pipeline can be found in Appendix B.3.
VideoSAM Model
VideoSAM extends the SAM architecture for HSV PD segmentation by integrating CNN-based initial segmentation with transformer-based refinement. The model utilizes a flexible prompt encoder, image encoder, and mask decoder to achieve pixel-level precision. Training and inference pipelines are optimized for multimodal datasets, ensuring robust segmentation across experimental conditions. Appendix B.4 provides full implementation details, equations, and illustrations.
Uncertainty Quantification
To assess the reliability of PD segmentation outputs, we employ a rigorous uncertainty quantification (UQ) framework that evaluates discretization errors in pixel-based measurements. We estimate errors in metrics such as contact line density and dry area fraction through iterative simulations and weighted frequency analysis. The methodology, including theoretical formulations and algorithmic details, is discussed in Appendix B.6.
@@ 4 Results and Discussions
Comparative Analysis of Dry Area Fraction and Contact Line Density
The analysis of boiling metrics, shown in Figure 3, reveals clear trends in dry area fraction and contact line density as functions of heat flux under saturated pool boiling conditions. Both segmentation methods—U-Net CNN and adaptive thresholding—indicate an upward trend consistent with expected boiling dynamics where vapor coverage increases with rising heat flux. Both techniques yield similar results at lower heat flux levels; however, significant deviations appear as the heat flux exceeds 140 kW/m². The U-Net consistently reports higher dry area fractions and contact line densities, particularly between 140-158 kW/m², where smaller, closely packed vapor bubbles dominate the interface. This difference underscores the U-Net’s sensitivity to finer vapor structures, often overlooked by the thresholding technique. Such accuracy in detecting small-scale features is crucial for understanding the dynamic behavior of the liquid-vapor interface and improving the prediction of heat transfer performance. Detailed statistical analyses are provided in Appendix C.1.1 for a more granular comparison of the segmentation errors.
@@
3D Histogram Analysis of Heat Flux vs. Bubble Size Distribution
The right-side plots of Figure 3 illustrate the bivariate distribution of bubble sizes across varying heat flux levels, highlighting the comparative performance of U-Net CNN and thresholding. At low heat flux, the bubble size distribution skews toward smaller bubbles, with the U-Net effectively capturing a broad range of bubble sizes. As the heat flux increases, the histogram shifts toward larger bubbles, reflecting the onset of vigorous boiling. The U-Net segmentation highlights a smooth progression of bubble size growth, with a notable concentration of medium-sized bubbles at intermediate flux levels. In contrast, the thresholding method struggles to detect smaller bubbles, particularly at lower flux levels, resulting in an incomplete distribution skewed toward larger sizes. The ability of the U-Net CNN to resolve fine-scale details contributes to a more comprehensive understanding of bubble dynamics across heat flux regimes, which is critical for optimizing thermal management systems. Full histogram comparisons and additional heat flux levels are detailed in Appendix C.1.2.
Comparison of Bubble Size Distribution and Visualizations
Figure 9 compares bubble size distributions for frames extracted from videos 8 and 9. The raw camera outputs, displayed as grayscale images, provide a baseline for assessing segmentation performance. The U-Net results, color-coded by bubble size (blue for small, green for medium, yellow for large, and orange for the largest), exhibit a diverse size range with clear distinctions between individual bubbles, even in densely packed regions. This fine-grained segmentation confirms the U-Net’s robustness in capturing intricate bubble dynamics, particularly in active boiling phases where bubble coalescence and fragmentation occur simultaneously. In comparison, the thresholding method highlights larger bubbles effectively but often merges or overlooks smaller bubbles, reducing segmentation accuracy. This difference is particularly pronounced in regions with closely packed bubbles, where U-Net consistently outperforms thresholding. These results emphasize the value of integrating U-Net segmentation into boiling experiments requiring precise bubble size characterization. Detailed visualizations of additional frames are included in Appendix C.1.3
Validation with Ground Truth Annotations
To assess the accuracy of the U-Net CNN model, we compared its results against ground truth annotations manually provided by five independent domain experts (Figure 10). The expert-annotated dry area fraction and contact line density values clustered around 0.17 and 0.05, respectively, with minimal inter-user variability. The U-Net segmentation aligns closely with the ground truth, achieving deviations below 2% for dry area fraction and under 8% for contact line density. In contrast, the thresholding method underestimates both metrics, particularly in frames with smaller bubbles or complex bubble interactions. The observed underestimation arises from thresholding’s limitations in detecting fine-scale details, leading to gaps in bubble identification. Error quantification (detailed in Appendix C.1.4) confirms that U-Net segmentation provides a more reliable approximation of ground-truth metrics, underscoring its suitability for high-precision boiling analysis.
Statistical Comparison of Segmentation Across Modalities
The robustness of the U-Net CNN model was further validated through a statistical comparison of boiling metrics across argon, nitrogen, and FC-72 modalities, as illustrated in Figure 12. In argon and nitrogen, both the U-Net and thresholding methods yielded comparable results, with dry area fractions ranging between 0.46-0.48 and contact line densities between 0.06-0.07. This consistency highlights the segmentation methods’ efficacy in relatively simpler boiling regimes. However, the FC-72 dataset revealed significant discrepancies between the two methods. The U-Net results showed a more clustered distribution of dry area fraction and contact line density, while thresholding exhibited greater variability and a higher incidence of outliers. These deviations align with FC-72’s complex boiling patterns, characterized by smaller, densely packed bubbles that challenge the thresholding technique. Probability density functions, cumulative distribution functions, and box plots (included in Appendix C.1.6) further confirm the U-Net’s superior performance in complex boiling scenarios, solidifying its role as a robust segmentation tool.
Qualitative and Quantitative Performance across Modalities
High-pressure water boiling experiments (10 bar, 3000 kW/m², and 40 bar, 3400 kW/m²) were used to evaluate the transferability and adaptability of the U-Net CNN model (Figure 11). Initially trained on liquid nitrogen data, the U-Net exhibited limitations when applied to water boiling, particularly in detecting large bubbles where hollow centers and false positives appeared. Fine-tuning the model with annotated high-pressure water frames significantly improved segmentation accuracy, as indicated by a close match between the fine-tuned U-Net results and expert-verified ground truth perimeters. These findings underscore the importance of domain-specific tuning for accurate segmentation in varying experimental conditions. This iterative approach—combining pre-trained models with fine-tuning—demonstrates the U-Net’s adaptability for diverse boiling environments, providing a foundation for robust analysis in high-pressure applications. Further results showing the qualitative perimeter visualization results across various modalities are provided in Appendix C.1.5. Meanwhile, the multi-modal quantitative results are presented here in Appendix C.1.7.
Uncertainty Quantification in Performance Parameters
To analyze the impact of discretization on boiling metrics, we investigated the effects of grid resolution and bubble size on dry area fraction and contact line density. These parameters were evaluated using numerical experiments, where grid resolutions and bubble radii were systematically varied. The methodology, detailed in Appendix C.2, compares the discretized measurements of bubble perimeter and area with their theoretical values, quantified using Percentage Relative Error (PRE) and Mean Error (ME) metrics. Figure 4 illustrates that perimeter measurements are particularly sensitive to grid size, transitioning from overestimation at fine resolutions to underestimation at coarser ones. In contrast, area measurements exhibit greater resilience, stabilizing more effectively across bubble radii and grid sizes.
Experimental Data Validation
A representative frame from the liquid argon dataset was selected and analyzed at a resolution of 12.6 $\mu$m/px under saturated boiling conditions to validate these numerical findings. The results, presented in Appendix  C.2.2, show the distributions of bubble perimeter, area, and radius (Figure 16). The probability distributions confirm that smaller bubbles dominate the sample, while a logarithmic scale highlights the presence of occasional larger bubbles. Convergence tests for PRE, performed over iteration milestones (5K to 20K), demonstrate that area calculations require more iterations to stabilize than the perimeter. At 20K iterations, numerical errors converge to negligible levels, balancing accuracy with computational efficiency.
@@
Grid Resolution and Bubble Size Influence
The impact of grid resolution and bubble size on measurement accuracy was further analyzed in Appendix  C.2.3. Figure 4 demonstrates the trends in ME and PRE for perimeter and area calculations. The results highlight a non-linear relationship between grid resolution and bubble characteristics, with perimeter errors exhibiting a distinct transition from overestimation to underestimation. Area errors remain predominantly negative, with minimal sensitivity to grid changes. These findings emphasize the need for fine grid resolutions to minimize the mistakes in perimeter estimation, particularly for smaller bubbles.
Erosion and Dilation Effects on Boundary Accuracy
We also examined the effects of boundary modifications, such as erosion and dilation, on measurement uncertainties. Appendix  C.2.4, and Figure 17 visually illustrate these effects. Dilation reduces edge-related artifacts, decreasing perimeter PRE from 13.8% to 6.6% (Table 9). This smoothing effect ensures closer alignment between perimeter estimates and ground truth values. While area PRE remains stable, the findings underscore the importance of selecting appropriate boundary conditions in digital image processing to refine bubble segmentation accuracy.
Uncertainty Quantification Across Modalities
The same uncertainty analysis was extended to FC-72, liquid argon (LAr), liquid nitrogen (LN2), and high-pressure water datasets, as detailed in Appendix  C.2.5. Table 2 summarizes the results, showing consistent PRE values for area measurements across all modalities. However, perimeter uncertainties vary significantly, with LN2 exhibiting the highest PRE (-8.4%) due to pronounced boundary irregularities. These discrepancies highlight the greater sensitivity of perimeter calculations to segmentation accuracy and pixel resolution, particularly in datasets with complex boiling dynamics.
@@
Modality $\alpha$ $\beta$ $\gamma$
@@
Symbol Definitions:
$\alpha$: Weighted Avg PRE Area (%)
$\beta$: Weighted Avg PRE Perimeter (%)
$\gamma$: Weighted Avg ME Perimeter ($\mu$m)
VideoSAM Segmentation Across Modalities
We evaluated the zero-shot generalization performance of VideoSAM across nitrogen, FC-72, and water datasets. Figure 18, detailed in Appendix  C.3.2, compares VideoSAM’s results with SAM and ground truth masks. VideoSAM demonstrates robust performance in nitrogen and FC-72 datasets, effectively capturing intricate bubble boundaries and outperforming SAM in terms of accuracy, IoU, and Dice scores. For instance, in the nitrogen dataset, VideoSAM achieves an IoU of 0.8317, compared to SAM’s 0.6702. However, its performance deteriorates in the water dataset, where low object variability and background noise undermine segmentation accuracy. These results highlight VideoSAM’s strength in handling dynamic environments with high bubble density and complex structures.
Multimodal Segmentation Performance
To further benchmark VideoSAM against U-Net and SAM, segmentation accuracy was compared across water, FC-72, nitrogen, and argon datasets. Table 3 and Figure 18 (Appendix  C.3.2) present the results for IoU and F1 scores. VideoSAM consistently outperforms other models in complex environments like nitrogen and FC-72, achieving IoU values of 0.83 and 0.80, respectively. In contrast, U-Net performs best in simpler datasets, such as water, with an IoU of 0.5619, underscoring its adaptability to low-contrast, uniform backgrounds. These results confirm the trade-offs between model generalization capacity and specialization, with VideoSAM excelling in dense, irregular boiling structures and U-Net delivering superior results for uniform datasets.